OpenStudy (anonymous):

Simplify the problem below. -5 × 9 --- ---- -21 + 6 Select one of the options below as your answer: A. -3 B. -5/3 C.5/3 D. 3

sam (.sam.):

\[\text{Is \it}\frac{-5\times9}{-21+6}~~?\]

OpenStudy (anonymous):

yes

sam (.sam.):

\[\frac{-5\times9}{-21+6}=\frac{-45}{-15}=3\]